OpenStudy (anonymous):

If sin theta=3/5 and theta is in quad 2 the exact form of sin (theta+ pi/6) is .....? literally have not figured out this question at all..

OpenStudy (anonymous):

I dont understand what you are asking, what do you mean by 'quad 2'.

OpenStudy (anonymous):

14mdaz, you have different quadrants when dealing with the unit circle and that is what he/she is referring to

OpenStudy (anonymous):

oh quadrants

OpenStudy (anonymous):

quadrant 2 is between pi/2 and pi

OpenStudy (anonymous):

What I don't get is what he / she is wanting from the exact form of sin (theta+ pi/6) is

OpenStudy (anonymous):

it will not be the cleanest value

OpenStudy (anonymous):

Is he wanting to add sin theta=3/5 + pi/6???

OpenStudy (anonymous):

no he wants theta plus pi/6 all in rads im guessing

OpenStudy (anonymous):

applied to sin function

OpenStudy (anonymous):

This whole question is just confusing me on top of that right now lol its asking for it to be in exact form

OpenStudy (anonymous):

Are you allowed to use a calculator or do you need to find \( \sin \theta = 3/5 \) without a calculator?

OpenStudy (anonymous):

Without a calculator sadly...

OpenStudy (anonymous):

hmm...

OpenStudy (anonymous):



OpenStudy (anonymous):

:3

OpenStudy (anonymous):

I think 14mdaz is explaining so I am going to step back

OpenStudy (anonymous):

uhh no i just doodled, its a highly inaccurate diagram

OpenStudy (anonymous):

I don't even think he knows

zepdrix (zepdrix):

Still need help on this one shorty? :)

OpenStudy (anonymous):

Yes pleaseee!

zepdrix (zepdrix):

You need to apply your Angle Sum Formula:\[\large\rm \sin(\alpha+\beta)=\sin \alpha \cos \beta+\sin \beta \cos \alpha\]

zepdrix (zepdrix):

Let's apply that before we do anything else

zepdrix (zepdrix):

\[\large\rm \sin\left(\theta+\frac{\pi}{6}\right)=\sin \theta \cos \frac{\pi}{6}+\sin\frac{\pi}{6}\cos \theta\]Do you understand how I applied that formula? :)

OpenStudy (anonymous):

Yeah I think I have done it that way and got the answer [3*sqrt(3) - 4]/10 but I don't think its right :/

zepdrix (zepdrix):

Oooo yes good job! That looks correct!! :)

OpenStudy (anonymous):

But how would the work look like?

OpenStudy (anonymous):

But when I go to check my answer something comes out differently.

OpenStudy (anonymous):

sin(theta)cos(pi/6) + cos(theta)sin(pi/6) = (3/5)(sqrt(3)/2) - (4/5)(1/2)

OpenStudy (anonymous):

The work is correct right?

zepdrix (zepdrix):

\[\large\rm \sin\left(\theta+\frac{\pi}{6}\right)=\left(\sin \theta\right)\left(\cos \frac{\pi}{6}\right)+\left(\sin\frac{\pi}{6}\right)\left(\cos \theta\right)\]\[\large\rm \sin\left(\theta+\frac{\pi}{6}\right)=\left(\frac{3}{5}\right)\left(\frac{\sqrt3}{2}\right)+\left(\frac{1}{2}\right)\left(\frac{-4}{5}\right)\]

zepdrix (zepdrix):

Plugging in the pieces :) ya looks right

OpenStudy (anonymous):

But is there a way to check it?

zepdrix (zepdrix):

Hmmm.

OpenStudy (anonymous):

He is using the Sum and Difference formula.

zepdrix (zepdrix):

\[\large\rm \sin(\theta)=\frac{3}{5}\qquad\to\qquad \sin^{-1}\frac{3}{5}=\theta\approx0.6435\] So then the sine of that angle theta... plus pi/6 should approximately give us (3sqrt3-4)/10, whatever decimal that works out to. Kind of a tough problem to check your work on :)

OpenStudy (anonymous):

You can easily check with a calculator. Just plugin the values

OpenStudy (anonymous):

Oh alright cuz my professor was all like you can check your work which is why I was asking